# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ D.1 Simulation studies
We first present additional simulation results in Table D.1, similar to those in Table 1, to analyze and compare further computational strategies related to the Pólya mixing parameter. Instead of independent Metropolis-Hastings, we employ a particle Gibbs sampler for sampling $\lambda$ with $N=10$ particles, combined with a blocked Gibbs sampler. Table D.1 reports the mixing behaviour and sampling efficiency of algorithms with a particle Gibbs sampler, depending on the utilization of an adaptive Pólya proposal scheme and three different strengths of dependence in data generation scenarios. Here, the acceptance rate for the particle Gibbs sampler is defined as the number of iterations when one of the new $N$ candidates is accepted as a new sample, divided by the total number of iterations.
Table D.1 shows that employing a particle Gibbs sampler for sampling $\lambda$ improves the effective sample size of $\lambda$, but at the cost of decreased sampling efficiency measured by effective sample size divided by seconds. Utilizing the adaptive Pólya proposal with a particle Gibbs sampler shows less significant impact compared to the adaptive Pólya proposal with independent Metropolis-Hastings, as both aim to improve the mixing behaviour of $\lambda$. These results suggest that the particle Gibbs sampler could enhance the mixing behaviour, which is particularly beneficial when the acceptance rate from independent Metropolis-Hastings is low. While this approach may improve mixing, it does not guarantee enhanced sampling efficiency.
Table D.2: Simulation results comparing latent logistic-beta process and Gaussian copula models with latent logistic-beta process data generation scenarios. Entries shown are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. LBP, logistic-beta process; RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
@@
Latent LBP, $\varrho=0.1$ Latent LBP 11.59 (0.15) 12.17 (0.20) 6.31 (0.10) 6.62 (0.12) 21.27 (0.24)
Gauss. copula 11.66 (0.15) 12.19 (0.20) 6.35 (0.09) 6.65 (0.12) 0.48 (0.01)
Latent LBP, $\varrho=0.2$ Latent LBP 8.54 (0.18) 8.73 (0.19) 4.67 (0.11) 4.77 (0.11) 17.87 (0.16)
Gauss. copula 8.59 (0.17) 8.76 (0.18) 4.70 (0.10) 4.79 (0.11) 0.44 (0.01)
Latent LBP, $\varrho=0.4$ Latent LBP 6.12 (0.19) 6.16 (0.19) 3.41 (0.11) 3.43 (0.11) 17.48 (0.16)
Gauss. copula 6.15 (0.17) 6.19 (0.18) 3.43 (0.10) 3.44 (0.11) 0.47 (0.01)
Table D.2 provides simulation results that compare latent logistic-beta process and Gaussian copula models that are similar to Table 2 but with data generation based on the latent logistic-beta process model. Compared to Table 2, where the Gaussian copula model achieved slightly better predictive performance, Table D.2 shows that the latent logistic-beta process model can achieve slightly better predictive performance of of latent success probabilities when the model is correctly specified. In terms of sampling efficiency, the latent logistic-beta process clearly outperforms the Gaussian copula model.
Table D.3: Additional simulation results comparing latent logistic-beta process and Gaussian copula models. Entries shown are average over 100 replicates, with Monte Carlo standard error in parentheses. LBP, logistic-beta process; MAE, mean absolute error of inferred & predicted success probabilities on training & test data; mean ESS/sec, Univariate effective sample sizes of latent success probabilities divided by running time in seconds, averaged across training dataset.
@@
Gaussian copula, $\varrho=0.1$ Latent LBP 9.34 (0.13) 9.67 (0.15) 9.44 (0.16) 73.44 (0.48)
Gaussian copula 9.24 (0.12) 9.60 (0.14) 0.49 (0.01) 2135.52 (44.88)
Gaussian copula, $\varrho=0.2$ Latent LBP 6.79 (0.14) 6.90 (0.14) 7.36 (0.15) 89.86 (0.38)
Gaussian copula 6.74 (0.14) 6.86 (0.15) 0.41 (0.01) 2882.76 (255.82)
Gaussian copula, $\varrho=0.4$ Latent LBP 4.82 (0.14) 4.84 (0.14) 7.45 (0.21) 89.98 (0.34)
Gaussian copula 4.80 (0.14) 4.83 (0.14) 0.47 (0.01) 2320.49 (84.65)
Latent LBP, $\varrho=0.1$ Latent LBP 9.03 (0.14) 9.54 (0.17) 9.31 (0.20) 73.50 (0.46)
Gauss. copula 9.10 (0.13) 9.58 (0.17) 0.48 (0.01) 2182.61 (47.37)
Latent LBP, $\varrho=0.2$ Latent LBP 6.64 (0.15) 6.78 (0.16) 7.36 (0.18) 88.65 (0.48)
Gauss. copula 6.67 (0.15) 6.81 (0.15) 0.45 (0.01) 2356.76 (50.61)
Latent LBP, $\varrho=0.4$ Latent LBP 4.84 (0.15) 4.86 (0.16) 7.53 (0.21) 89.91 (0.46)
Gauss. copula 4.85 (0.14) 4.88 (0.15) 0.48 (0.01) 2305.82 (85.17)
@@
We provide additional summaries in Table D.3 that compare latent logistic-beta process and Gaussian copula models. Similar to Table 2, Table D.3 shows that the in-sample and predictive performance of the latent logistic-beta process model in terms of mean absolute error is comparable to the Gaussian copula model, showing the flexibility of logistic-beta process despite model misspecification. When data are generated from the latent logistic-beta process model, the in-sample and predictive performance of the latent logistic-beta process model is slightly better than the Gaussian copula model. We also present sampling efficiency in terms of the average of univariate effective sample size and algorithm running times. Logistic-beta process models have a clear computational advantage, likely due to the conditionally conjugate updating scheme with an efficient blocked Gibbs sampling algorithm.
D.2 Real data analysis
@@
We divide the collaborative perinatal project data available from the BNPmix package into two groups based on maternal smoking status during pregnancy, resulting in data sizes of $n=1023$ (smoking) and $n=1290$ (nonsmoking), respectively. We centred and scaled the data before the analysis, and we set the upper bound on the number of components as $H=20$ for both the logistic-beta dependent Dirichlet process and the logit stick-breaking process. We considered semiconjugate priors for mixture component parameters in expression (13) as $(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(0_{2},I_{2}),\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(1,1)$, $h=1,\dots,H$ for both models.
@@
We run Algorithm 2 for logistic-beta dependent Dirichlet process, and Algorithm 1 of Rigon and Durante, (2021) for logit stick-breaking process to carry out the posterior inference, both implemented with R for comparison purposes. We ran a total of 35,000 Markov chain Monte Carlo iterations for both models, and the first 5,000 samples were discarded as burn-in, executed on an Intel(R) Xeon(R) Gold 6132 CPU with a 96GB memory environment. The analysis of trace plots shows satisfactory convergence for both models. The running time of Markov chain Monte Carlo samplers under settings 1, 2, and 3 are approximately 5.4, 7.5, and 10.2 minutes for the logistic-beta dependent Dirichlet process and 4.9, 5.6, and 6.2 minutes for the logit stick-breaking process, respectively, for the analysis of smoking group data. This shows the computational tractability of the logistic-beta dependent Dirichlet process model is similar to the logit stick-breaking process model, and we note that substantial time improvements are possible with lower-level programming languages.
The choice of basis functions for logistic-beta dependent Dirichlet process and logit stick-breaking process models with covariate-dependent weights are similar to Rigon and Durante, (2021). We consider a natural cubic spline basis with $q=6$ degrees of freedom as unnormalized feature maps (including the intercept term so that splines are not restricted to meet at origin), obtained from R package splines with automatically chosen knot locations. Basis functions are normalized to have unit Euclidean norm. The left panel of Figure D.1 presents unnormalized and normalized basis functions on the DDE exposure data for the smoking group.
Figure D.1: Comparison of prior properties based on unnormalized (dashed) and normalized feature map (solid). (Left) Unnormalized and normalized basis functions with natural cubic splines for the smoking group. (Right) Comparison of prior co-clustering probabilities and prior expected number of clusters $E(K_{1023})$ given a DDE level under logit stick-breaking process with prior on basis coefficients $\alpha\sim N_{6}(0_{6},\pi^{2}/3I_{6})$.
Under the logit stick-breaking process with unnormalized basis functions $\varphi(x)$ and prior on coefficients $\alpha\sim N_{q}(0_{q},\sigma_{\alpha}^{2}I_{q})$, we have $\eta(x)=\varphi(x)^{\mathrm{\scriptscriptstyle T}}\alpha\sim N(0,\sigma_{\alpha}^{2}\|\varphi(x)\|_{2}^{2}$), whose variance varies over $x$. The corresponding stick-breaking ratio $\sigma\{\eta(x)\}$ follows a logit-normal distribution $F_{V(x)}$ with parameters 00 and $\sigma_{\alpha}^{2}\|\varphi(x)\|_{2}^{2}$, so that prior co-clustering probability $\mu_{2}(x)/(2\mu_{1}(x)-\mu_{2}(x))$ also varies with $x$ where $\mu_{1}(x)=E(V(x))$ and $\mu_{2}=E(V(x)^{2})$ for $V(x)\sim F_{V(x)}$; see Lemma 1 of Dunson and Park, (2008). This is shown as a dashed line in the top right panel of Figure D.1 when $\sigma^{2}_{\alpha}=\pi^{2}/3$ (setting 2), showing marginal prior properties varies with $x$. This is true following previous default choices (Rigon and Durante,, 2021; Horiguchi et al.,, 2022; Kang and Kottas,, 2022). In contrast, under logit stick-breaking process with normalized basis functions $\phi(x)$ and prior $\alpha\sim N_{q}(0_{q},\sigma_{\alpha}^{2}I)$, we have $\eta(x)=\phi(x)^{\mathrm{\scriptscriptstyle T}}\alpha\sim N_{1}(0,\sigma_{\alpha}^{2})$ that does not depend on $x$. The corresponding prior co-clustering probability $\mu_{2}(x)/(2\mu_{1}(x)-\mu_{2}(x))$ also does not depend on $x$, shown as a solid line of the top right panel of Figure D.1 when $\sigma^{2}_{\alpha}=\pi^{2}/3$ (Setting 2). The same fact also holds for the prior expected number of clusters in $n$ samples, denoted as $E(K_{n})$. The bottom right panel of Figure D.1 shows $E(K_{1023})$ under logit stick-breaking process with unnormalized basis (dashed) and normalized basis (solid) functions, obtained from Monte Carlo simulations.
Based on these prior properties, we use normalized basis functions for both the logistic-beta dependent Dirichlet process and the logit stick-breaking process and set their hyperparameters to match their prior co-clustering probabilities. For setting 1, we set $b=0.2$ for the logistic-beta dependent Dirichlet process and $\sigma_{\alpha}^{2}=10^{2}$ for the logit stick-breaking process, which leads to prior co-clustering probabilities $5/6=0.833$ (logistic-beta dependent Dirichlet process) and $0.854$ (logit stick-breaking process), respectively. For setting 2, we set $b=1$ for the logistic-beta dependent Dirichlet process and $\sigma_{\alpha}^{2}=\pi^{2}/3$ for the logit stick-breaking process, where $\pi^{2}/3$ corresponds to the variance of the standard logistic distribution. This leads to prior co-clustering probabilities corresponding to $1/2=0.5$ (logistic-beta dependent Dirichlet process) and $0.515$ (logit stick-breaking process), respectively. For setting 3, we set $b=2$ for the logistic-beta dependent Dirichlet process, $\sigma_{\alpha}^{2}=0.2^{2}$ for the logit stick-breaking process, which leads to prior co-clustering probabilities corresponding to $1/3=0.333$ (logistic-beta dependent Dirichlet process) and $0.338$ (logit stick-breaking process), respectively.
Figure D.2 presents the estimated preterm birth probabilities given DDE exposures for the nonsmoking group of size $n=1290$. Similar to Fig. 5, the estimated preterm birth probabilities show an overall increasing pattern as the DDE exposure level rises. However, unlike the logistic-beta dependent Dirichlet process, the hyperparameter choice of the logit stick-breaking process model has a substantial impact on uncertainty estimates.
@@
Figure D.2: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the logistic-beta dependent Dirichlet process model (top) and the logit stick-breaking process model (bottom) with three different hyperparameter settings. DDE exposure data are represented by small vertical bars.

### v3

@@ Acknowledgements
    We first present additional simulation results in Table S.3.1, similar to those in Table 1, to analyze and compare the computational strategies involved. First, we compare with a naive version of the Gibbs sampler where Step 2 of Algorithm 9 is replaced with sampling $\lambda$ from full conditional $\pi(\lambda\mid\bm{\omega},\bm{\eta},\mathbf{z})$, referred to as “Non-blocked”. Also, instead of independent Metropolis-Hastings, we employ a particle Gibbs sampler for sampling $\lambda$ with $N=10$, combined with a blocked Gibbs sampler.
    Table S.3.1 reports the mixing behavior and sampling efficiency of $\lambda$ of algorithms under different computational strategies. Here, the acceptance rate for the particle Gibbs sampler is defined as the number of iterations when one of the new $N$ candidates is accepted as a new sample, divided by the total number of iterations. The blocking scheme has a substantial impact on the mixing behavior, leading to an effective sample size increase of up to 50 times. Also, it shows that employing a particle Gibbs sampler improves the effective sample size of $\lambda$, but at the cost of decreased sampling efficiency measured by effective sample size divided by seconds. Utilizing the adaptive Pólya proposal with a particle Gibbs sampler shows less significant impact compared to the adaptive Pólya proposal with independent Metropolis-Hastings, as both aim to improve the mixing behavior of $\lambda$. These results highlight the benefits of the blocking strategy, which is made possible due to the normal mixture representation of logistic-beta. It also suggests that the particle Gibbs sampler could enhance the mixing behavior, which is particularly beneficial when the acceptance rate from independent Metropolis-Hastings is low. While this approach may improve mixing, it does not guarantee enhanced sampling efficiency.
    Table S.3.2: Nonparametric binary regression simulation results comparing latent logistic-beta process (LBP) and Gaussian copula models (GauCop). Entries are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
@@
    LBP, $\varrho=0.1$ LBP 11.59 (0.15) 12.17 (0.20) 6.31 (0.10) 6.62 (0.12) 21.27 (0.24)
    GauCop 11.66 (0.15) 12.19 (0.20) 6.35 (0.09) 6.65 (0.12) 0.48 (0.01)
    LBP, $\varrho=0.2$ LBP 8.54 (0.18) 8.73 (0.19) 4.67 (0.11) 4.77 (0.11) 17.87 (0.16)
    GauCop 8.59 (0.17) 8.76 (0.18) 4.70 (0.10) 4.79 (0.11) 0.44 (0.01)
    LBP, $\varrho=0.4$ LBP 6.12 (0.19) 6.16 (0.19) 3.41 (0.11) 3.43 (0.11) 17.48 (0.16)
    GauCop 6.15 (0.17) 6.19 (0.18) 3.43 (0.10) 3.44 (0.11) 0.47 (0.01)
    Table S.3.2 provides simulation results that compare latent logistic-beta process and Gaussian copula models that are similar to Table 2 but with data generation based on the latent logistic-beta process model. Compared to Table 2, where the Gaussian copula model achieved slightly better predictive performance, Table S.3.2 shows that the latent logistic-beta process model can achieve slightly better predictive performance of of latent success probabilities when the model is correctly specified. In terms of sampling efficiency, the latent logistic-beta process clearly outperforms the Gaussian copula model.
    Table S.3.3: Additional simulation results comparing latent logistic-beta process (LBP) and Gaussian copula models (GauCop). Entries shown are average over 100 replicates, with Monte Carlo standard error in parentheses. MAE, mean absolute error of inferred & predicted success probabilities on training & test data; mean ESS/sec, Univariate effective sample sizes of latent success probabilities divided by running time in seconds, averaged across training dataset.
@@
    GauCop, $\varrho=0.1$ LBP 9.34 (0.13) 9.67 (0.15) 9.44 (0.16) 73.44 (0.48)
    GauCop 9.24 (0.12) 9.60 (0.14) 0.49 (0.01) 2135.52 (44.88)
    GauCop, $\varrho=0.2$ LBP 6.79 (0.14) 6.90 (0.14) 7.36 (0.15) 89.86 (0.38)
    GauCop 6.74 (0.14) 6.86 (0.15) 0.41 (0.01) 2882.76 (255.82)
    GauCop, $\varrho=0.4$ LBP 4.82 (0.14) 4.84 (0.14) 7.45 (0.21) 89.98 (0.34)
    GauCop 4.80 (0.14) 4.83 (0.14) 0.47 (0.01) 2320.49 (84.65)
    LBP, $\varrho=0.1$ LBP 9.03 (0.14) 9.54 (0.17) 9.31 (0.20) 73.50 (0.46)
    GauCop 9.10 (0.13) 9.58 (0.17) 0.48 (0.01) 2182.61 (47.37)
    LBP, $\varrho=0.2$ LBP 6.64 (0.15) 6.78 (0.16) 7.36 (0.18) 88.65 (0.48)
    GauCop 6.67 (0.15) 6.81 (0.15) 0.45 (0.01) 2356.76 (50.61)
    LBP, $\varrho=0.4$ LBP 4.84 (0.15) 4.86 (0.16) 7.53 (0.21) 89.91 (0.46)
    GauCop 4.85 (0.14) 4.88 (0.15) 0.48 (0.01) 2305.82 (85.17)
@@
    We provide additional summaries in Table S.3.3 that compare latent logistic-beta process and Gaussian copula models. Similar to Table 2, Table S.3.3 shows that the in-sample and predictive performance of the latent logistic-beta process model in terms of mean absolute error is comparable to the Gaussian copula model, showing the flexibility of logistic-beta process despite model misspecification. When data are generated from the latent logistic-beta process model, the in-sample and predictive performance of the latent logistic-beta process model is slightly better than the Gaussian copula model. We also present sampling efficiency in terms of the average of univariate effective sample size and algorithm running times. Logistic-beta process models have a clear computational advantage, likely due to the conditionally conjugate updating scheme with an efficient blocked Gibbs sampling algorithm.
    S.3.2 Bayesian density regression simulation studies
    First, we provide a detailed description of the regression errors and density errors we used to evaluate the simulation results, where we closely followed Wade and Inacio, (2025). We consider $n_{\mathrm{test}}=100$ test covariates $x^{*}_{1},\dots,x^{*}_{100}$ equally spaced over the domain, $[0,1]$ for scenario A and $[-2,10]$ for scenario B. Also, we consider the grid of responses $\{y_{g}\}$ of size $500$, equally spaced over $[-1,2]$ for scenario A and $[-1,10]$ for scenario B. The regression errors are calculated as
    Regression $\text{Regression error}=\left[\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\{m(y\mid x_{i}^{*})-\hat{m}(y\mid x_{i}^{*})\}^{2}\right]^{1/2}$
    where we used the posterior mean estimate for $\hat{m}(y\mid x_{i}^{*})$. Also, the density errors are calculated as
    $\text{Density error}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\sum_{g=1}^{500}|p(y_{g}\mid x_{i}^{*})-\hat{p}(y_{g}\mid x_{i}^{*})|(y_{2}-y_{1})$
    where we used the posterior mean estimate for $\hat{p}(y_{g}\mid x_{i}^{*})$.
    We set the upper bound on the number of components as $H=20$ for all models. For the priors on atom processes, we used normal prior on regression coefficients $(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(\bm{0}_{2},10^{2}\mathbf{I}_{2})$ and gamma prior on precision $\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(1,1)$ for $h=1,\dots,H$. All algorithms are executed on Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU. For all three models, we run a total of 25,000 Markov chain Monte Carlo iterations, and the first 20,000 samples are discarded as burn-in.
    Figure S.3.1: Density regression simulation results under two different data generation scenarios with sample size $n=2000$ and three different hyperparameter settings based on prior tie probability. Density error and regression error quantify the accuracy of conditional density and conditional mean estimates averaged across test covariates, respectively.
    Figure S.3.1 provides density regression simulation results that are similar to Figure 5 but with sample size $n=2000$. Compared to Figure 5, the performance of the linear DDP mixture model is not improved in both scenarios despite the increased sample size. This highlights the benefits of having covariate-dependent mixture weights, where the linear DDP mixture model is not flexible enough to accommodate the complexities. The LB-DDP and LSBP mixture models show improved performance, but similar to Figure 5, the LSBP mixture model is also sensitive to the choice of hyperparameter for the setting when the prior tie probability is around 0.34.
    S.3.3 Pregnancy outcome data analysis
@@
    We divide the collaborative perinatal project data available from the BNPmix package into two groups based on maternal smoking status during pregnancy, resulting in data sizes of $n=1023$ (smoking) and $n=1290$ (nonsmoking), respectively. We centred and scaled the data before the analysis, and we set the upper bound on the number of components as $H=20$ for both the logistic-beta dependent Dirichlet process and the logit stick-breaking process. We considered semiconjugate priors for mixture component parameters in expression (15) as $(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(\bm{0}_{2},\mathbf{I}_{2}),\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(1,1)$, $h=1,\dots,H$ for both models.
@@
    We run Algorithm 2 for logistic-beta dependent Dirichlet process, and Algorithm 1 of Rigon and Durante, (2021) for logit stick-breaking process to carry out the posterior inference, both implemented with R for time comparison purposes. We ran a total of 35,000 Markov chain Monte Carlo iterations for both models, and the first 5,000 samples were discarded as burn-in, executed on an Intel(R) Xeon(R) Gold 6132 CPU with a 96GB memory environment. The analysis of trace plots shows satisfactory convergence for both models. The running time of Markov chain Monte Carlo samplers under settings 1, 2, and 3 are approximately 5.4, 7.5, and 10.2 minutes for the logistic-beta dependent Dirichlet process and 4.9, 5.6, and 6.2 minutes for the logit stick-breaking process, respectively, for the analysis of smoking group data. This shows the computational tractability of the logistic-beta dependent Dirichlet process model is similar to the logit stick-breaking process model, and we note that substantial time improvements are possible with lower-level programming languages.
@@
    Figure S.3.2: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the logistic-beta dependent Dirichlet process model (top) and the logit stick-breaking process model (bottom) with three different hyperparameter settings corresponding to prior tie probabilities approximately 0.84, 0.50, and 0.34. DDE exposure data are represented by small vertical bars.
    Figure S.3.2 presents the estimated preterm birth probabilities given DDE exposures for the non-smoking group of size $n=1290$. Similar to Figure 6, the estimated preterm birth probabilities show an overall increasing pattern as the DDE exposure level rises. However, unlike the logistic-beta dependent Dirichlet process, the hyperparameter choice of the logit stick-breaking process model has a substantial impact on uncertainty estimates.